# Linear-Cost Vecchia Approximation of Multivariate Normal Probabilities

### v2

@@ 6.4 Stress test for VMET under constant correlation
Even under a most challenging covariance structure for the Vecchia approximation, VMET still provided a better trade-off between estimation errors and computation cost than all methods except for MET, which achieved higher accuracy with substantially more computation cost. VMET with multi-level MC used $m_{2}/2=m_{1}=m$ and $10N_{2}=N_{1}=N$, which, despite higher accuracy than VMET at the same $m$, led to a less favorable trade-off between accuracy and computation cost. It is worth mentioning that the constant correlation structure is rarely used in practice while those in Table 1 are more representative, especially among spatial applications. Therefore, Figure 5 serves as a ‘stress test’ for the proposed VMET method, providing numerical support for its performance under general non-spatial covariance structures, while Figures 3 and 4 are more indicative on its performance in various applications.
@@ 7 Analysis of Partially Censored Data
$\mathbf{x}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}),\qquad z_{i}=\begin{cases}x_{i},&\mbox{if }x_{i}>b_{i},\\ \mbox{NA},&\mbox{otherwise},\end{cases}$ (26)
@@
$f(\mathbf{z})=f_{\mathbf{x}_{1:n_{1}}}(\mathbf{z}_{1:n_{1}})\Pr(\mathbf{x}_{n_{1}+1:n}<\mathbf{b}_{n_{1}+1:n}\mid\mathbf{x}_{1:n_{1}}=\mathbf{z}_{1:n_{1}}),$ (27)
@@
Notice that $\mathbf{x}_{c(i)}$ for $i>n_{1}$ may include observed $x_{j}$ (i.e., with $j\leq n_{1}$). We can view (28) as a limit case of the general MVN probability in (2) with $\mathbf{a}_{1:n_{1}}\rightarrow\mathbf{z}_{1:n_{1}}\leftarrow\mathbf{b}_{1:n_{1}}$ and $\mathbf{a}_{n_{1}+1:n}=-\bm{\infty}$.
@@ 8 Conclusions
Similar to other scalable methods (e.g., Genton et al., 2018; Cao et al., 2021; Nascimento & Shaby, 2022), our proposed VMET method amounts to an approximation to the covariance in the original MVN probability, and hence estimation bias is unavoidable. However, the Vecchia estimation bias is typically negligible relative to the MC errors for small values of the conditioning-set size $m$ (e.g., $30$ to $50$), making the bias from the Vecchia approximation a worthwhile trade-off for much higher computation scalability. Furthermore, the Vecchia approximation used in VMET provides parallel-computation capability with respect to the number of responses $n$, in addition to the parallelization with respect to the sample size $N$. One future research avenue is studying other families of proposal densities. Generally speaking, the design of proposal densities should consider how to 1) draw samples from the proposal density and 2) optimize the density function parameters. The convex-concave property, described in Proposition 4, for other families of proposal densities may no longer hold, necessitating innovative criteria for optimizing the proposal density, similar to (11). This paper only considers MVN probabilities with positive-definite covariance matrices; a direction for future research is to study Vecchia approximation of MVN probabilities subject to linear constraints, $\mbox{Pr}(\mathbf{a}\leq\mathbf{G}\mathbf{x}\leq\mathbf{b})$, where $\mathbf{G}$ can be rectangular, resulting in rank-deficient covariance matrices.

### v3

@@ 6.4 Stress test for VMET under constant correlation
Even under a most challenging covariance structure for the Vecchia approximation, VMET still provided a better trade-off between estimation errors and computation cost than all methods except for MET, which achieved higher accuracy with substantially more computation cost. VMET with multi-level MC used $m_{2}/2=m_{1}=m$ and $10N_{2}=N_{1}=N$, which, despite higher accuracy than VMET at the same $m$, led to a less favorable trade-off between accuracy and computation cost. It is worth mentioning that the constant correlation structure is rarely used in practice while those in Table 1 are more representative, especially among spatial applications. Therefore, Figure 4 serves as a ‘stress test’ for the proposed VMET method, providing numerical support for its performance under general non-spatial covariance structures, while Figures 2 and 3 are more indicative of its performance in various applications.
@@ 7 Analysis of Partially Censored Data
$\mathbf{x}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}),\qquad\text{ and }z_{i}=x_{i}\ \mbox{if }x_{i}>b_{i}\mbox{ and }\mbox{NA}\ \mbox{otherwise},$
@@
$f(\mathbf{z})=f_{\mathbf{x}_{1:n_{1}}}(\mathbf{z}_{1:n_{1}})\Pr(\mathbf{x}_{n_{1}+1:n}<\mathbf{b}_{n_{1}+1:n}\mid\mathbf{x}_{1:n_{1}}=\mathbf{z}_{1:n_{1}}),$
@@
Notice that $\mathbf{x}_{c(i)}$ for $i>n_{1}$ may include observed $x_{j}$ (i.e., with $j\leq n_{1}$). We can view (21) as a limit case of (1) with $\mathbf{a}_{1:n_{1}}\rightarrow\mathbf{z}_{1:n_{1}}\leftarrow\mathbf{b}_{1:n_{1}}$ and $\mathbf{a}_{n_{1}+1:n}=-\bm{\infty}$.
@@ 8 Conclusions
Similar to other scalable methods (e.g., Genton et al., 2018; Cao et al., 2021; Nascimento & Shaby, 2022), our proposed VMET method amounts to an approximation to the covariance in the original MVN probability, and hence estimation bias is unavoidable. However, the Vecchia estimation bias is typically negligible relative to the MC errors for small values of the conditioning-set size $m$ (e.g., $30$ to $50$), making the bias from the Vecchia approximation a worthwhile trade-off for much higher computation scalability. Furthermore, the Vecchia approximation used in VMET provides parallel-computation capability with respect to the number of responses $n$, in addition to the parallelization with respect to the sample size $N$. One future research avenue is studying other families of proposal densities. Generally speaking, the design of proposal densities should consider how to 1) draw samples from the proposal density and 2) optimize the density function parameters. The convex-concave property, described in Proposition 4, for other families of proposal densities may no longer hold, necessitating innovative criteria for optimizing the proposal density, similar to (10). This paper only considers MVN and MVT probabilities with positive-definite covariance matrices; a direction for future research is to study Vecchia approximation of MVN probabilities subject to linear constraints, $\mbox{Pr}(\mathbf{a}\leq\mathbf{G}\mathbf{x}\leq\mathbf{b})$, where $\mathbf{G}$ can be rectangular, resulting in rank-deficient covariance matrices.